OpenStudy (anonymous):

Factor 2x^2 - 11x - 21 A. (2x + 3)(x - 7) B. (x + 3)(2x - 7) C. (2x - 3)(x + 7) D. (2x + 7)(x - 3) Please explain how to do this, don't just give the answer. Thanks.

OpenStudy (anonymous):

once my friend is avalable ill ask her to answer an exsplain this too you ok..=)

OpenStudy (anonymous):

can i give you ans

OpenStudy (anonymous):

ans is a

OpenStudy (anonymous):

(ax+b)(cx+d) = acx^2 + (bc+ad)x + bd that's the general rule. your problem has -21 = bd, and all your choices have -21 when you multiply the, so that does not eliminate any. your problem has 2x^2 and all your choices have ac=2, so that does not eliminate any of them, either your problem has -11x, so (bc+ad) = -11, and we have to check to see which option has this a. (2x+3)(x-7) a=2, b=3, c=1, d=-7 bc+ad= 3-14 = -11 bingo! just what we wanted. so @akshay1234 is right. b. (x+3)(2x-7) a=1, b=3, c=2, d=-7, bc+ad = 6-7=-1 not -11 try c. and d. but they are not right, as a. is the correct answer.